OpenStudy (anonymous):

Whats the highest common factor of 4 and 10

myininaya (myininaya):

4/2=2 10/2=5 no more common factors after you divide by 2 so gcf=2

OpenStudy (anonymous):

oh yeah i forgot how to work it out thank you :P

myininaya (myininaya):

np you could write it like this write prime factorizations for both 4=2*2 10=2*5 they only have a two in common so gcf=2 --------------- example:12 and 18 12=2*2*3 18=2*3*3 gcf=2*3